OpenStudy (anonymous):

Did I get the right answer? What is the equation of the quadratic graph with a focus of (5, 6) and a directrix of y = 2? I got f(x)=1/8(x+5)^2+4

OpenStudy (amistre64):

vertex shares the same x part ... so -5 is better just at a cursory view

OpenStudy (amistre64):

+4 is fine since its the middle of 6 and 2

OpenStudy (tkhunny):

8 is fine, since 6-4 = 2 and 4-2 = 2 and 2*4 = 8

OpenStudy (amistre64):

and the 1/8 looks fine with the 1/(4p)

OpenStudy (anonymous):

Can you explain why the 5 turns into a negative? That confuses me :P

OpenStudy (amistre64):

spose we have the function: y = x^2 we can move this thing left or right by adjusting the value of the x part y = (x+h)^2 the adjustment moves the x=0 to something else, x+h = 0 when x=-h but a -x is not a convention that is used soo x-h gets played into action

OpenStudy (amistre64):

spose we have a zero when x=5 (5+h) = 0 when h=-5 (x + (-5))

OpenStudy (anonymous):

Okay so if the 5 were a negative, would it turn into a positive?

OpenStudy (amistre64):

of course: say we have a zero when x=-5 (-5+h) = 0 when h=5 (x+(+5))

OpenStudy (anonymous):

Ohh okay, thank-you!

OpenStudy (amistre64):

another view of it x+h = 0 when h=-x

OpenStudy (amistre64):

good luck :)